Shadow:

These should be more your speed @Vocaloid Give me a moment to write them out on here then solve them.

Shadow:

\[X = [3, 5, 0, 2]\] a. \[\sum X \] b. \[\sum X^2\] c. \[\sum X + 1\] d. \[\sum (X + 1)\]

Shadow:

a. \[\sum X = \sum (3, 5, 0, 2) = 3 + 5 + 0 + 2 = 10\] b. \[\sum X^2 = \sum (3, 5, 0, 2)^2 = 3^2 + 5^2 + 0^2 + 2^2 = 9 + 25 + 0 + 4 = 38\] c. \[\sum X + 1 = \sum (3, 5, 0, 2) +1 = (3 + 5 + 0 + 2)+ 1 = 11\] d. \[\sum (X + 1) = \sum (3 + 1, 5 + 1, 0+1, 2 +1) = 4 + 6 + 1 + 3 = 14\]

Shadow:

Next set, \[X = (3, 2, 5, 1, 3)\] a. \[\sum X^2 = \sum (3^2, 2^2, 5^2, 1^2, 3^2) = 9 + 4 + 25 + 1 + 9 = 48\] b. \[(\sum X)^2 = \sum(3, 2, 5, 1, 3)^2 = (3+2+5+1+3)^2 = 14^2 = 196\] c. \[\sum (X - 1) = (3 -1, 2-1, 5-1, 1-1, 3-1) = 2 + 1 + 4 + 0 + 2 = 9\] d. \[\sum (X - 1)^2 = \sum(3 -1, 2-1, 5-1, 1-1, 3-1)^2 = (2 + 1 + 4 + 0 + 2)^2 = 9^2 = 81\]

Shadow:

am like sleep deprived so making sure I don't fail the first hw ._.

Vocaloid:

everything looks good to me